# A Causally Informed Pretraining Approach for Multimodal Foundation Models: Applications in Remote Sensing

### v3

@@ 10. Discussion and Conclusion
In this paper, we proposed a novel pretraining task for building foundation models, which we call Causally Informed Variable Step Forecasting (CI-VSF). Using remote sensing as a case study, we showed that the embeddings produced by our approach capture causal dependencies between weather and satellite data (i.e captures the impact of weather on the land cover). These embeddings result in improved fine-tuned models for downstream tasks compared to the embeddings produced by traditional pretraining methods. Specifically, we observed significant performance gains in soil moisture-related tasks (soil moisture prediction and soil moisture forecasting), pixel-wise crop mapping, and spectral imagery-based tasks (missing image prediction and future image forecasting).
Note that the weather is not the sole determining factor for the land cover, as it can be impacted by human actions (e.g.,urbanization) or other natural factors not directly covered by weather (e.g., damage to crops via pest infestation). Such downstream tasks will not benefit from our approach. That said, there are numerous other applications where our methodology could have a significant impact. Some examples include: Wildfire tracking and progression forecasting (Wildfire spread is influenced by wind speed and land dryness, both of which are determined by weather conditions), Crop yield forecasting (Weather patterns directly impact crop growth, and can help mitigate potential losses), Snow cover melt forecasting (snowmelt behavior is based on temperature and precipitation patterns can aid water resource management).
Future work will include the development of a full-scale foundation model on a scale much larger than used in this study with only 50,000 samples, and make it available for use by the community in a wide range of geoscience applications.
@@ A.1. Pretraining Stages
A.2. Pretraining Results
@@
Figure 12. Image Forecast Downstream Task comparison. Row 1 depicts a crop field where, Green arrows depict regions of growth, and Red arrows depict regions of harvest, CI-VSF captures both these phenomena better than SM-VSF. Row 2 depicts a case where CI-VSF adds snowfall accurately compared to SM-VSF. Row 3 depicts a case where CI-VSF does not change land cover due to terrain but SM-VSF adds false greenness. Larger version available in Appendix Figure
A.3. Soil Mositure Forecasting
Region of Analysis:Figure 13 depicts the geographic locations of the 6 tiles used in the soil moisture experiments.
Architecture and Implementation details: To capture the past context soil moisture values in the architecture, we use a unidirectional LSTM and add the corresponding hidden state embeddings with their counterparts in the $Emb_{STW}$ (We follow the same strategy of how we combine weather and satellite imagery embeddings). This series will then be used to forecast a value in the future based on the forecast day for that sample (delta T). Since the output is a single value for each sample time series, we use a series of linear layers with Relu on the final embdeding to lower the dimension to a single value. Since we are finetuning, we freeze the weights of the encoder and update only the weights of the decoder and the soil moisture LSTM for 50 epochs using Mean Absolute Error Loss and AdamW Optimizer with a learning rate of 0.0001.
A.4. Soil Moisture Prediction
Architecture and Implementation details Since our output is a single value for each timestamp in the series, we use a series of shared linear layers with Relu following each Embedding in the series to get to our output at each timestamp. For each sample, since the images were sampled at random temporally across the entire year the resulting satellite image series has uneven time intervals. This leads us to also pass the day of year series as input as well. All models were finetuned for 70 epochs. Like the previous task, the encoder weights were frozen and only the shared linear heads were finetuned using Mean Absolute Error Loss and AdamW optimizer with a learning rate of 0.0001.
@@ A.4. Soil Moisture Prediction
Crop Class SM-MR MM-MR SM-VSF CI-VSF
Corn 0.4135 0.4717 0.4489 0.5708
Cotton 0.8346 0.8403 0.9055 0.9125
Winter Wheat 0.1156 0.0770 0.1043 0.1777
Tomatoes 0.7680 0.7637 0.7223 0.7341
Grapes 0.7398 0.7502 0.7447 0.7543
Almonds 0.3851 0.4386 0.2073 0.2990
Walnut 0.0238 0.1417 0.4494 0.5384
Pistachio 0.5070 0.6734 0.6200 0.7003
Alfalfa 0.6892 0.7271 0.7072 0.7057
Grass 0.7760 0.8436 0.7715 0.8445
Urban 0.6111 0.6408 0.6229 0.6191
Average 0.5331 0.5789 0.5731 0.6233
A.5. Pixel Wise Crop Mapping
Region of Analysis:Figure 8 depicts the Cropland Data Layer(CDL) labels and the region of analysis (Sentinel Tile T11SKA) for our crop mapping downstream task. As can be seen our region of analysis lies in the heart of the California Central Valley and contains numerous crop classes.
Architecture and Implementation details: To perform crop mapping we would need an architecture that gives us a pixel wise output, in particular, we would need a decoder that takes the embedding series given by the foundation model encoder, i.e $Emb_{STW}$, using it to construct a pixel wise classification map. To map this embedding series to a pixel wise map, we follow an attention based approach, similar to WSTATT(Ravirathinam et al., 2024). This strategy assigns a weightage to each timestamp and does an aggregated sum to form a multitemporal attention-based embedding. This multitemporal embedding is then acted on by a series of upscaling and convolution layers with activation, and an output Linear layer to form a pixel wise map. Figure 9 depicts the general layout of the architecture used for the crop mapping downstream task. The embedding series mentioned would correspond to the series $Emb_{STW}$, i.e the output series of the encoder. In all methods(CI-VSF, SM-MR, etc.) the encoder weights are fixed from the respective pretraining task and only the attention mechanism and decoder layers are finetuned. All models finetuning is done with only 2018 data for 40 epochs using Cross Entropy Loss and Adam optimizer with a learning rate of 0.0001.
Performance: Classwise F1 score performance on the test set can be seen in Table 6. We can see a good improvement in classes where weather influences crop growth directly, such as Corn and Walnut.
A.6. Missing Image Prediction
@@
A.7. Future Image Forecasting
Architecture and Implementation details: Since we are forecasting imagery in the future, we can keep the overall architecture the same as during the pretraining(figure 4 for the CI-VSF variant) but reinitialise the decoder weights. We used a input series of length 6, kept the encoder fixed and only finetune the decoder layers for just 10 epochs for both the SM-VSF and CI-VSF variants using Mean Square Error Loss on the final image using the Adam Optimiser and learning rate of 0.0001. We also did not use masking during the finetuning stage.
Performance Larger version of comparison of forecasted images downstream task is present in Figure 12

### v4

@@ 13. Discussion and Conclusion
In this paper, we introduced Knowledge-Guided Variable-Step Forecasting (KG-VSF), a new pre-training objective for multimodal foundation models that explicitly leverages causal relationships between modalities. Using remote sensing as a case study, we demonstrated that KG-VSF embeddings capture the causal influence of weather (drivers) on land-surface conditions observed via satellite imagery (responses). Compared to traditional pre-training methods such as masked reconstruction and variable-step forecasting, KG-VSF consistently produced embeddings that yielded superior fine-tuned performance across multiple downstream applications, including soil moisture prediction and forecasting, pixel-wise crop type mapping, and spectral imagery-based tasks such as missing-image reconstruction and future-image forecasting.
Beyond these examples, the proposed framework opens new possibilities for a wide range of applications where understanding causal interactions among variables is critical. In geoscience, these include wildfire progression forecasting, crop yield prediction, and snowmelt estimation, all of which depend on how environmental drivers influence system responses. More broadly, the same paradigm can be applied to other domains—such as healthcare, where patient outcomes depend on complex interactions among clinical and external factors, or finance, where macroeconomic indicators causally affect market dynamics. By embedding causal reasoning directly into pre-training, KG-VSF provides a principled foundation for developing models that are more physically consistent, interpretable, and transferable across scientific and societal domains.
Looking ahead, we see the main opportunity not merely in scaling up model size, but in expanding the conceptual and methodological space of knowledge-guided pre-training by developing new objectives, architectures, and evaluation frameworks that incorporate domain knowledge in diverse forms. While building full-scale geoscience foundation models will ultimately require substantial computational resources, our results suggest that even modest-scale implementations of knowledge-guided pre-training can yield significant benefits. Given sufficient computational capacity, the KG-VSF paradigm could enable large-scale foundation models that capture causal relationships across diverse domains and data modalities
@@ A.1. Pretraining Stages and Benefit
A.1.1. Single-phase vs Two-phase approach for pre-training KGVSF
In multimodal contexts like ours, where weather acts as the driver variable and satellite imagery as the response, directly combining masked reconstruction and forecasting objectives in a single loss often can lead to optimization imbalance. The reconstruction objective, being spatially dense and easier to minimize, would tend to dominate the joint loss early in training. As a result, the encoder focuses on learning local texture and spatial smoothness priors from the imagery rather than exploring conditional dependencies across modalities. This yields embeddings that are spatially rich but causally shallow. The forecasting component, which depends on cross-modal reasoning (e.g., inferring how rainfall or temperature anomalies manifest in land cover patterns weeks later), receives weaker gradients during early training, preventing effective modeling of conditional dynamics.
Our two-phase strategy directly addresses this issue. In Phase 1 (masked reconstruction), each modality’s encoder (satellite and weather) and the multimodal sequence encoder are first trained to capture robust representations of spatial and temporal regularities within their respective modalities. This establishes a stable representational foundation that is unbiased toward either input channel. In Phase 2 (forecasting), we introduce a forecaster that must predict future satellite imagery conditioned on both historical satellite context and weather up to the forecast date. Because the reconstruction task is already well learned, the optimization in this phase focuses on aligning weather-driven temporal transitions with satellite responses, thereby embedding causal relationships in the shared latent space.
In summary, while single-phase multi-objective pretraining may be simpler, it tends to emphasize easily optimizable spatial reconstruction signals at the expense of cross-modal causal learning. The proposed two-phase procedure, though slightly more complex, explicitly guides the model from learning modality-specific structure to capturing conditional, knowledge-guided relationships—an essential step for producing generalizable and causally meaningful geospatial embeddings. To empirically show the benefits of the two phase approach, we compare embeddings from both these models on the pixel wise crop mapping task and soil moisture prediction. From Table 8, we can see that in both cases, the two phase approach outperformed the single phase multiobjective approach, showing that empirically as well, the two phase approach shows benefit.
Table 8. Empirical Comparison of the two phase vs single phase pretraining approaches
Crop Mapping Soil Moisture
Avg F1 Score R Square
Single Phase Multi objective Pretraining 0.7189 0.7221
Two Phase wise Pretraining 0.7602 0.8081
@@ A.1.1. Single-phase vs Two-phase approach for pre-training KGVSF
A.2. Masking
Figure 12 shows an example of such masking for a image series of 4 4x4 grids with 50% masking. From the Figure, we can see that in each timestamp image there are 8 patches masked and focusing on a particular patch location along the temporal dimension we notice that 2 patches are available.
A.2.1. Masking Ratio Comparison
We conduct an experiment to figure out the optimal masking ratio. We train multiple models using 60k samples on different masking ratios and finetuned for the downstream tasks of pixel wise crop mapping and soil moisture estimation.
Table 9. Comparison of performance on downstream tasks of crop mapping and soil moisture prediction from KGVSF models pretrained with different masking ratios
Crop Mapping Soil Moisture
Masking Ratio used in Pretraining KGVSF Avg F1 Score R Square
25% 0.7391 0.7340
50% 0.7602 0.8081
75% 0.7197 0.7689
From Table 9, we can see that a moderate amount of masking leads to the best results. This results aligns with the masking results from other papers such as Satmae(cong2022satmae) and ScaleMAE(reed2023scalemaescaleawaremaskedautoencoder), where they also reported best results with a moderate amount of masking.
A.3. Pretraining Results
@@
A.4. Pixel Wise Crop Mapping
A.4.1. Architecture and Implementation details
: To perform crop mapping we would need an architecture that gives us a pixel wise output, in particular, we would need a decoder that takes the embedding series given by the pretrained model encoder, i.e $Emb_{STW}$, using it to construct a pixel wise classification map. To map this embedding series to a pixel wise map, we follow an attention based approach, similar to WSTATT(wstatt). This strategy assigns a weightage to each timestamp and does an aggregated sum to form a multitemporal attention-based embedding. This multitemporal embedding is then acted on by a series of upscaling and convolution layers with activation, and an output Linear layer to form a pixel wise map. Figure 10 depicts the general layout of the architecture used for the crop mapping downstream task. The embedding series mentioned would correspond to the series $Emb_{STW}$, i.e the output series of the encoder. In all methods(KG-VSF, SM-MR, etc.) the encoder weights are fixed from the respective pretraining task and only the attention mechanism and decoder layers are finetuned. All models finetuning is done with only 2018 data for 40 epochs using Cross Entropy Loss and Adam optimizer with a learning rate of 0.0001.
A.4.2. Global Dataset Results
Our data for finetuning comes from Sentinel2 and ERA5 land data for the region of the T11SKA Sentinel tile in the California Central Valley, a region rich in various crop classes and has been used for crop type mapping in various other works(statt; calcrop21; wstatt) from the years 2018 and 2019. Figure 9 depicts the Cropland Data Layer(CDL) labels and the region of analysis (Sentinel Tile T11SKA) for our crop mapping downstream task. As can be seen our region of analysis lies in the heart of the California Central Valley and contains numerous crop classes. Like other works, we get our labels for this region from the Cropland Data Layer(CDL), an annually released land cover map for the entire continuous US by the USDA. A diagram of the CDL labels and geographic location of T11SKA tile can be seen in Appendix Figure 9.
In this problem setting, we use 10 spectral images (biweekly from May to Sept) and weather data (within that timeframe) to get embeddings from each framework and use that embedding series to predict pixel wise crop labels at a 10m resolution. Similar to WSTATT (wstatt), we adopt a grid based training method, by splitting the entire region into train, validation and test grids and also follow their preprocessing steps including combination and erosion. Due to the prediction related nature of this task, models from all 4 pretraining tasks are finetuned. A decoding head for semantic segmentation is appended and finetuned using cross entropy loss. Pretrained Encoder to obtain embeddings is kept fixed.
Table 10 compares the performance of KG-VSF and other pretraining approaches when finetuned for crop mapping task. Though finetuning is done using 2018 data, testing is done on 2019 data, assessing the robustness of the approaches. We can see that KG-VSF performs better than other variants. We also observe that MM-MR performs better than SM-MR, which is because inclusion of weather in input is in general better for crop mapping task (wstatt), however MM-MR and MM-VSF lack capturing the relationship between modalities, thus its inferior performance when compared to KG-VSF. Finally, the ability of KG-VSF’s embeddings to generalize across years after finetuning indicates that the embeddings have captured crucial information that go beyond what is just present in spectral imagery. Classwise F1 score performance on the test set can be seen in Table 12. We can see a good improvement in classes where weather influences crop growth directly, such as Corn and Walnut.
2019 Test 11 class average F1 Scores
SM-MR $M$-MR SM-VSF $M$-VSF KG-VSF
Average 0.5331 0.5789 0.5731 0.5767 0.6233
A.4.3. Comparison of Performance across different timeframes
From the Crop mapping downstream task, we saw we used 10 Sentinel-2 timestamps focused around the growing season of crops (i.e May to to get our results. To ensure that this is the optimal number of we experimented with multiple temporal variants to assess the model’s sensitivity to the number of observation samples. As shown in Table 11, we can see that as long as the input data is focused on the growing season, the performance remains good.
Table 11. Comparison on downstream task of crop mapping across different input timeframes
Crop Mapping
Input Timeframe Avg F1 Score
Growing season(May to Sept) sparse (monthly) 0.7417
Growing season(May to Sept) all timestamps 0.7237
All year data sparse (monthly) 0.6734
All year data all timestamps ($\sim$40–60 time steps/year) 0.6993
Growing season(May to Sept) frequent (biweekly -10 time steps) (Our choice) 0.7602
@@ A.4.3. Comparison of Performance across different timeframes
Crop Class SM-MR $M$-MR SM-VSF $M$-VSF KG-VSF
Corn 0.4135 0.4717 0.4489 0.4836 0.5708
Cotton 0.8346 0.8403 0.9055 0.9037 0.9125
Winter Wheat 0.1156 0.0770 0.1043 0.1129 0.1777
Tomatoes 0.7680 0.7637 0.7223 0.7172 0.7341
Grapes 0.7398 0.7502 0.7447 0.7355 0.7543
Almonds 0.3851 0.4386 0.2073 0.2509 0.2990
Walnut 0.0238 0.1417 0.4494 0.4053 0.5384
Pistachio 0.5070 0.6734 0.6200 0.6245 0.7003
Alfalfa 0.6892 0.7271 0.7072 0.6973 0.7057
Grass 0.7760 0.8436 0.7715 0.7831 0.8445
Urban 0.6111 0.6408 0.6229 0.6301 0.6191
Average 0.5331 0.5789 0.5731 0.5767 0.6233
A.5. Soil Moisture Prediction
A.5.1. Architecture and Implementation details
Since our output is a single value for each timestamp in the series, we use a series of shared linear layers with Relu following each Embedding in the series to get to our output at each timestamp. For each sample, since the images were sampled at random temporally across the entire year the resulting satellite image series has uneven time intervals. This leads us to also pass the day of year series as input as well. All models were finetuned for 70 epochs. Like the previous task, the encoder weights were frozen and only the shared linear heads were finetuned using Mean Absolute Error Loss and AdamW optimizer with a learning rate of 0.0001.
A.5.2. Global Dataset Results
Figure 14 depicts the geographic locations of the 6 tiles used in the soil moisture experiments. From the same 6 Sentinel tiles used in the soil moisture prediction problem, we obtain the Sentinel-2 satellite data and ERA5 weather data to serve as our input modalities and the ’soil_moisture_am’ band from SMAP to serve as our groundtruth for soil moisture(li2021improved). For sample generation, like the previous section, each tile is split into 100 grids and a 60-20-20 train-val-test split is done, and after preprocessing, we retain 1800 samples for training (300 from each tile), 600 for validation (100 from each tile), and 600 for testing (100 from each tile). Due to the prediction related nature of this task, models from all 4 pretraining tasks are finetuned. A decoding head for prediction is appended and finetuned using Mean Absolute Error loss. Pretrained Encoder to obtain embeddings for past data is kept fixed.
The performance of finetuned models on this task are presented in Table 13. The first experiment conducted was training using samples from all tiles and testing on samples from the test grids of each tile. From this in region sample testing, we can see KG-VSF’s finetuned model outperforms other models quite significantly. To explore robustness of learned representations, we also explore cross region testing. In this setting, during finetuning, samples from 1 tile are hidden and during testing, models are tested samples from that hidden tile. For example, models are trained on the 1500 samples from 5 tiles (T10SEJ, T14SKC, T15TUH, T16SBF, T14RQT) and tested on the 100 test samples from the hidden tile (T11SKA). We run this setting hiding each tile one at a time, shown in the lower portion of Table 13. From these results, KG-VSF’s model is able to perform cross region soil moisture estimation better than other models in all tiles, showing that embeddings from KG-VSF are able to capture causality and retain this information even when tested on regions not included in finetuning. This highlights how Knowledge Guided pretraining strategies help in downstream tasks.
Soil Moisture Prediction Finetuned Models
TEST Tile SM-MR $M$-MR SM-VSF $M$-VSF KG-VSF
In region Testing: Train on 6 tiles, Test on all 6 tiles
Mean Absolute Error Reported(Lower the better)
All 0.0615 0.0458 0.0483 0.0411 0.0282
Cross-Region Testing: Train on 5 tiles, Test on the 6th.
Mean Absolute Error Reported(Lower the better)
T11SKA 0.1113 0.0847 0.1121 0.0913 0.0695
T15TUH 0.1283 0.1365 0.1181 0.1132 0.0834
T14SKC 0.1159 0.1275 0.1312 0.1093 0.0958
T16SBF 0.0821 0.0631 0.0895 0.0619 0.0544
T10SEJ 0.1003 0.0718 0.1011 0.0742 0.0587
T14RQT 0.0815 0.0579 0.0658 0.0571 0.0558
A.6. Soil Mositure Forecasting
Architecture and Implementation details: To capture future weather we use a unidirectional LSTM and add the corresponding hidden state embeddings with the final embedding in the $Emb_{STW}$ series. This multimodal embedding is then used to predict the soil moisture at a future date. Since the output is a single value for each sample time series, we use a series of linear layers with Relu on the final embdeding to lower the dimension to a single value. Since we are finetuning, we freeze the weights of the pretrained encoder and update only the weights of the appended decoder layers and the weather LSTM for 50 epochs using Mean Absolute Error Loss and AdamW Optimizer with a learning rate of 0.0001.
Global Dataset Results Figure 14 depicts the geographic locations of the 6 tiles used in the soil moisture experiments. Here too, To estimate a soil moisture value in the future, we provide past satellite image time series and weather data to get embeddings from each pretraining objective. Then using those embeddings and weather till a future date we estimate the soil moisture at a future date. We use an input timeframe of 6 satellite images and predict the soil moisture value at a random 7th timestamp in the future. Each tile is split into 100 grids and a 60-20-20 train-val-test split is done, ensuring no overlap in regions. Multiple samples are generated from each grid, and after preprocessing, we retain 1800 samples for training, 400 for validation, and 400 for testing.
From Table 14, in short forecast ranges (i.e less than 25 days) most finetuned models are able to forecast soil moisture quite well with a moderate difference between them, however KG-VSF is top performing. However, when we move to larger forecast ranges, we can see a big jump in the performance of KG-VSF finetuned model compared to other finetuned models. Even though, as the forecast range increases, errors of all models go up, the errors by the other model go up significantly higher compared to KG-VSF’s model. This can be attributed to the fact that the embeddings from the KG-VSF model are able to capture the causal relationships between weather satellite imagery better than the other counterparts, leading to a performance boost.
Table 14. Comparison of models fine-tuned from different pretraining tasks on soil moisture forecasting downstream task. Table reports average R Squared Score of forecasted SMAP values for different forecast day ranges. In the test region soil moisture ranged from 0.062 to 0.4998 with a mean of 0.2134 and standard deviation of 0.0997
Soil Moisture Forecasting Downstream Task
(R Squared Scores Reported (higher the better))
Future Forecast Day Range SM-MR $M$-MR SM-VSF $M$-VSF KG-VSF
0 - 25 days 0.7231 0.8436 0.7349 0.8607 0.9501
25 - 50 days 0.7098 0.7724 0.7403 0.8019 0.9332
50 - 100 days 0.5818 0.7619 0.6114 0.7735 0.9216
More than 100 days 0.5136 0.7593 0.5864 0.7653 0.9012
A.7. Missing Image Prediction
@@
A.8. Future Image Forecasting
Architecture and Implementation details: Since we are forecasting imagery in the future, we can keep the overall architecture the same as during the pretraining(figure 4 for the KG-VSF variant) and load the pretrained weights and freeze them. We used a input series of length 6 and did not use masking during this downstream task.